# How Fair is Your Diffusion Recommender Model?

### v1

@@ Acknowledgements.
- Hoogeboom et al. (2022) Emiel Hoogeboom, Victor Garcia Satorras, Clément Vignac, and Max Welling. 2022. Equivariant Diffusion for Molecule Generation in 3D. In ICML (Proceedings of Machine Learning Research, Vol. 162). PMLR, 8867–8887.
- Jiang et al. (2024) Yangqin Jiang, Yuhao Yang, Lianghao Xia, and Chao Huang. 2024. DiffKG: Knowledge Graph Diffusion Model for Recommendation. In WSDM. ACM, 313–321.
- Li et al. (2022) Xiang Li, John Thickstun, Ishaan Gulrajani, Percy Liang, and Tatsunori B. Hashimoto. 2022. Diffusion-LM Improves Controllable Text Generation. In NeurIPS.
@@
- Li et al. (2023) Zihao Li, Aixin Sun, and Chenliang Li. 2023. DiffuRec: A Diffusion Model for Sequential Recommendation. CoRR abs/2304.00686 (2023).
@@
- Liang et al. (2018) Dawen Liang, Rahul G. Krishnan, Matthew D. Hoffman, and Tony Jebara. 2018. Variational Autoencoders for Collaborative Filtering. In WWW. ACM, 689–698.
@@
- Liu et al. (2023) Chengyi Liu, Wenqi Fan, Yunqing Liu, Jiatong Li, Hang Li, Hui Liu, Jiliang Tang, and Qing Li. 2023. Generative Diffusion Models on Graphs: Methods and Applications. In IJCAI. ijcai.org, 6702–6711.
@@
- Liu et al. (2021) Qigang Liu, Lifeng Mu, Vijayan Sugumaran, Chongren Wang, and Dongmei Han. 2021. Pair-wise ranking based preference learning for points-of-interest recommendation. Knowl. Based Syst. 225 (2021), 107069.
- Ma et al. (2024) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Zhanhui Kang. 2024. Plug-In Diffusion Model for Sequential Recommendation. In AAAI. AAAI Press, 8886–8894.
@@
- Mansoury et al. (2022) Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. 2022. A Graph-Based Approach for Mitigating Multi-Sided Exposure Bias in Recommender Systems. ACM Trans. Inf. Syst. 40, 2 (2022), 32:1–32:31.
@@
- Mao et al. (2021) Kelong Mao, Jieming Zhu, Xi Xiao, Biao Lu, Zhaowei Wang, and Xiuqiang He. 2021. UltraGCN: Ultra Simplification of Graph Convolutional Networks for Recommendation. In CIKM. ACM, 1253–1262.
@@
- Naghiaei et al. (2022) Mohammadmehdi Naghiaei, Hossein A. Rahmani, and Yashar Deldjoo. 2022. CPFair: Personalized Consumer and Producer Fairness Re-ranking for Recommender Systems. In SIGIR. ACM, 770–779.
@@
- Niu et al. (2020) Chenhao Niu, Yang Song, Jiaming Song, Shengjia Zhao, Aditya Grover, and Stefano Ermon. 2020. Permutation Invariant Graph Generation via Score-Based Generative Modeling. In AISTATS (Proceedings of Machine Learning Research, Vol. 108). PMLR, 4474–4484.
@@
- Qin et al. (2024) Yifang Qin, Hongjun Wu, Wei Ju, Xiao Luo, and Ming Zhang. 2024. A Diffusion Model for POI Recommendation. ACM Trans. Inf. Syst. 42, 2 (2024), 54:1–54:27.
@@
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian Personalized Ranking from Implicit Feedback. In UAI. AUAI Press, 452–461.
@@
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In CVPR. IEEE, 10674–10685.
@@
- Sarwar et al. (2001) Badrul Munir Sarwar, George Karypis, Joseph A. Konstan, and John Riedl. 2001. Item-based collaborative filtering recommendation algorithms. In WWW. ACM, 285–295.
@@
- Shen et al. (2023) Xudong Shen, Chao Du, Tianyu Pang, Min Lin, Yongkang Wong, and Mohan S. Kankanhalli. 2023. Finetuning Text-to-Image Diffusion Models for Fairness. CoRR abs/2311.07604 (2023).
@@
- Shenbin et al. (2020) Ilya Shenbin, Anton Alekseev, Elena Tutubalina, Valentin Malykh, and Sergey I. Nikolenko. 2020. RecVAE: A New Variational Autoencoder for Top-N Recommendations with Implicit Feedback. In WSDM. ACM, 528–536.
- Steck (2019) Harald Steck. 2019. Embarrassingly Shallow Autoencoders for Sparse Data. In WWW. ACM, 3251–3257.
@@
- Sun et al. (2020) Jianing Sun, Wei Guo, Dengcheng Zhang, Yingxue Zhang, Florence Regol, Yaochen Hu, Huifeng Guo, Ruiming Tang, Han Yuan, Xiuqiang He, and Mark Coates. 2020. A Framework for Recommending Accurate and Diverse Items Using Bayesian Graph Convolutional Neural Networks. In KDD. ACM, 2030–2039.
@@
- Togashi et al. (2024) Riku Togashi, Kenshi Abe, and Yuta Saito. 2024. Scalable and Provably Fair Exposure Control for Large-Scale Recommender Systems. In WWW. ACM, 3307–3318.
@@
- Vassøy and Langseth (2024) Bjørnar Vassøy and Helge Langseth. 2024. Consumer-side fairness in recommender systems: a systematic survey of methods and evaluation. Artif. Intell. Rev. 57, 4 (2024), 101.
@@
- Wang et al. (2017) Jun Wang, Lantao Yu, Weinan Zhang, Yu Gong, Yinghui Xu, Benyou Wang, Peng Zhang, and Dell Zhang. 2017. IRGAN: A Minimax Game for Unifying Generative and Discriminative Information Retrieval Models. In SIGIR. ACM, 515–524.
- Wang et al. (2023) Wenjie Wang, Yiyan Xu, Fuli Feng, Xinyu Lin, Xiangnan He, and Tat-Seng Chua. 2023. Diffusion Recommender Model. In SIGIR. ACM, 832–841.
@@
- Wu et al. (2023b) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, Hui Xiong, and Enhong Chen. 2023b. A Survey on Large Language Models for Recommendation. CoRR abs/2305.19860 (2023).
@@
- Wu et al. (2023a) Shiwen Wu, Fei Sun, Wentao Zhang, Xu Xie, and Bin Cui. 2023a. Graph Neural Networks in Recommender Systems: A Survey. ACM Comput. Surv. 55, 5 (2023), 97:1–97:37.
@@
- Xu et al. (2022) Minkai Xu, Lantao Yu, Yang Song, Chence Shi, Stefano Ermon, and Jian Tang. 2022. GeoDiff: A Geometric Diffusion Model for Molecular Conformation Generation. In ICLR. OpenReview.net.
@@
- Yang et al. (2016) Dingqi Yang, Daqing Zhang, and Bingqing Qu. 2016. Participatory Cultural Mapping Based on Collective Behavior Data in Location-Based Social Networks. ACM Trans. Intell. Syst. Technol. 7, 3 (2016), 30:1–30:23.
@@
- Yang et al. (2024) Ling Yang, Zhilong Zhang, Yang Song, Shenda Hong, Runsheng Xu, Yue Zhao, Wentao Zhang, Bin Cui, and Ming-Hsuan Yang. 2024. Diffusion Models: A Comprehensive Survey of Methods and Applications. ACM Comput. Surv. 56, 4 (2024), 105:1–105:39.
@@
- Yu et al. (2024) Junliang Yu, Xin Xia, Tong Chen, Lizhen Cui, Nguyen Quoc Viet Hung, and Hongzhi Yin. 2024. XSimGCL: Towards Extremely Simple Graph Contrastive Learning for Recommendation. IEEE Trans. Knowl. Data Eng. 36, 2 (2024), 913–926.
@@
- Zhao et al. (2022b) Minghao Zhao, Le Wu, Yile Liang, Lei Chen, Jian Zhang, Qilin Deng, Kai Wang, Xudong Shen, Tangjie Lv, and Runze Wu. 2022b. Investigating Accuracy-Novelty Performance for Graph-based Collaborative Filtering. In SIGIR. ACM, 50–59.
@@
- Zhao et al. (2022a) Wayne Xin Zhao, Yupeng Hou, Xingyu Pan, Chen Yang, Zeyu Zhang, Zihan Lin, Jingsen Zhang, Shuqing Bian, Jiakai Tang, Wenqi Sun, Yushuo Chen, Lanling Xu, Gaowei Zhang, Zhen Tian, Changxin Tian, Shanlei Mu, Xinyan Fan, Xu Chen, and Ji-Rong Wen. 2022a. RecBole 2.0: Towards a More Up-to-Date Recommendation Library. In CIKM. ACM, 4722–4726.

### v2

@@ Acknowledgements.
- Hoogeboom et al. (2022) Emiel Hoogeboom, Victor Garcia Satorras, Clément Vignac, and Max Welling. 2022. Equivariant Diffusion for Molecule Generation in 3D. In International Conference on Machine Learning, ICML 2022, 17-23 July 2022, Baltimore, Maryland, USA (Proceedings of Machine Learning Research, Vol. 162), Kamalika Chaudhuri, Stefanie Jegelka, Le Song, Csaba Szepesvári, Gang Niu, and Sivan Sabato (Eds.). PMLR, 8867–8887. https://proceedings.mlr.press/v162/hoogeboom22a.html
- Hou et al. (2024) Yu Hou, Jin-Duk Park, and Won-Yong Shin. 2024. Collaborative Filtering Based on Diffusion Models: Unveiling the Potential of High-Order Connectivity. In Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2024, Washington DC, USA, July 14-18, 2024, Grace Hui Yang, Hongning Wang, Sam Han, Claudia Hauff, Guido Zuccon, and Yi Zhang (Eds.). ACM, 1360–1369. https://doi.org/10.1145/3626772.3657742
- Jiang et al. (2024) Yangqin Jiang, Yuhao Yang, Lianghao Xia, and Chao Huang. 2024. DiffKG: Knowledge Graph Diffusion Model for Recommendation. In Proceedings of the 17th ACM International Conference on Web Search and Data Mining, WSDM 2024, Merida, Mexico, March 4-8, 2024, Luz Angelica Caudillo-Mata, Silvio Lattanzi, Andrés Muñoz Medina, Leman Akoglu, Aristides Gionis, and Sergei Vassilvitskii (Eds.). ACM, 313–321. https://doi.org/10.1145/3616855.3635850
- Kose and Shen (2024) Oyku Deniz Kose and Yanning Shen. 2024. FairWire: Fair Graph Generation. In Advances in Neural Information Processing Systems 38: Annual Conference on Neural Information Processing Systems 2024, NeurIPS 2024, Vancouver, BC, Canada, December 10 - 15, 2024, Amir Globersons, Lester Mackey, Danielle Belgrave, Angela Fan, Ulrich Paquet, Jakub M. Tomczak, and Cheng Zhang (Eds.). http://papers.nips.cc/paper_files/paper/2024/hash/e105d1cabb6ee2495595e2baf25493b5-Abstract-Conference.html
- Li et al. (2024a) Mufei Li, Eleonora Kreacic, Vamsi K. Potluru, and Pan Li. 2024a. GraphMaker: Can Diffusion Models Generate Large Attributed Graphs? Trans. Mach. Learn. Res. 2024 (2024). https://openreview.net/forum?id=0q4zjGMKoA
- Li et al. (2022) Xiang Lisa Li, John Thickstun, Ishaan Gulrajani, Percy Liang, and Tatsunori B. Hashimoto. 2022. Diffusion-LM Improves Controllable Text Generation. In Advances in Neural Information Processing Systems 35: Annual Conference on Neural Information Processing Systems 2022, NeurIPS 2022, New Orleans, LA, USA, November 28 - December 9, 2022, Sanmi Koyejo, S. Mohamed, A. Agarwal, Danielle Belgrave, K. Cho, and A. Oh (Eds.). http://papers.nips.cc/paper_files/paper/2022/hash/1be5bc25d50895ee656b8c2d9eb89d6a-Abstract-Conference.html
@@
- Li et al. (2024b) Zihao Li, Aixin Sun, and Chenliang Li. 2024b. DiffuRec: A Diffusion Model for Sequential Recommendation. ACM Trans. Inf. Syst. 42, 3 (2024), 66:1–66:28. https://doi.org/10.1145/3631116
@@
- Liang et al. (2018) Dawen Liang, Rahul G. Krishnan, Matthew D. Hoffman, and Tony Jebara. 2018. Variational Autoencoders for Collaborative Filtering. In Proceedings of the 2018 World Wide Web Conference on World Wide Web, WWW 2018, Lyon, France, April 23-27, 2018, Pierre-Antoine Champin, Fabien Gandon, Mounia Lalmas, and Panagiotis G. Ipeirotis (Eds.). ACM, 689–698. https://doi.org/10.1145/3178876.3186150
@@
- Liu et al. (2023) Chengyi Liu, Wenqi Fan, Yunqing Liu, Jiatong Li, Hang Li, Hui Liu, Jiliang Tang, and Qing Li. 2023. Generative Diffusion Models on Graphs: Methods and Applications. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence, IJCAI 2023, 19th-25th August 2023, Macao, SAR, China. ijcai.org, 6702–6711. https://doi.org/10.24963/IJCAI.2023/751
@@
- Liu et al. (2021) Qigang Liu, Lifeng Mu, Vijayan Sugumaran, Chongren Wang, and Dongmei Han. 2021. Pair-wise ranking based preference learning for points-of-interest recommendation. Knowl. Based Syst. 225 (2021), 107069. https://doi.org/10.1016/J.KNOSYS.2021.107069
- Ma et al. (2024) Haokai Ma, Ruobing Xie, Lei Meng, Xin Chen, Xu Zhang, Leyu Lin, and Zhanhui Kang. 2024. Plug-In Diffusion Model for Sequential Recommendation. In Thirty-Eighth AAAI Conference on Artificial Intelligence, AAAI 2024, Thirty-Sixth Conference on Innovative Applications of Artificial Intelligence, IAAI 2024, Fourteenth Symposium on Educational Advances in Artificial Intelligence, EAAI 2014, February 20-27, 2024, Vancouver, Canada, Michael J. Wooldridge, Jennifer G. Dy, and Sriraam Natarajan (Eds.). AAAI Press, 8886–8894. https://doi.org/10.1609/AAAI.V38I8.28736
@@
- Mansoury et al. (2022) Masoud Mansoury, Himan Abdollahpouri, Mykola Pechenizkiy, Bamshad Mobasher, and Robin Burke. 2022. A Graph-Based Approach for Mitigating Multi-Sided Exposure Bias in Recommender Systems. ACM Trans. Inf. Syst. 40, 2 (2022), 32:1–32:31. https://doi.org/10.1145/3470948
@@
- Mao et al. (2021) Kelong Mao, Jieming Zhu, Xi Xiao, Biao Lu, Zhaowei Wang, and Xiuqiang He. 2021. UltraGCN: Ultra Simplification of Graph Convolutional Networks for Recommendation. In CIKM ’21: The 30th ACM International Conference on Information and Knowledge Management, Virtual Event, Queensland, Australia, November 1 - 5, 2021, Gianluca Demartini, Guido Zuccon, J. Shane Culpepper, Zi Huang, and Hanghang Tong (Eds.). ACM, 1253–1262. https://doi.org/10.1145/3459637.3482291
- Medda et al. (2025) Giacomo Medda, Francesco Fabbri, Mirko Marras, Ludovico Boratto, and Gianni Fenu. 2025. GNNUERS: Fairness Explanation in GNNs for Recommendation via Counterfactual Reasoning. ACM Trans. Intell. Syst. Technol. 16, 1 (2025), 6:1–6:26. https://doi.org/10.1145/3655631
@@
- Naghiaei et al. (2022) Mohammadmehdi Naghiaei, Hossein A. Rahmani, and Yashar Deldjoo. 2022. CPFair: Personalized Consumer and Producer Fairness Re-ranking for Recommender Systems. In SIGIR ’22: The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, Madrid, Spain, July 11 - 15, 2022, Enrique Amigó, Pablo Castells, Julio Gonzalo, Ben Carterette, J. Shane Culpepper, and Gabriella Kazai (Eds.). ACM, 770–779. https://doi.org/10.1145/3477495.3531959
@@
- Niu et al. (2020) Chenhao Niu, Yang Song, Jiaming Song, Shengjia Zhao, Aditya Grover, and Stefano Ermon. 2020. Permutation Invariant Graph Generation via Score-Based Generative Modeling. In The 23rd International Conference on Artificial Intelligence and Statistics, AISTATS 2020, 26-28 August 2020, Online [Palermo, Sicily, Italy] (Proceedings of Machine Learning Research, Vol. 108), Silvia Chiappa and Roberto Calandra (Eds.). PMLR, 4474–4484. http://proceedings.mlr.press/v108/niu20a.html
- Purificato et al. (2022) Erasmo Purificato, Ludovico Boratto, and Ernesto William De Luca. 2022. Do Graph Neural Networks Build Fair User Models? Assessing Disparate Impact and Mistreatment in Behavioural User Profiling. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management, Atlanta, GA, USA, October 17-21, 2022, Mohammad Al Hasan and Li Xiong (Eds.). ACM, 4399–4403. https://doi.org/10.1145/3511808.3557584
- Purificato et al. (2024) Erasmo Purificato, Ludovico Boratto, and Ernesto William De Luca. 2024. Toward a Responsible Fairness Analysis: From Binary to Multiclass and Multigroup Assessment in Graph Neural Network-Based User Modeling Tasks. Minds Mach. 34, 3 (2024), 33. https://doi.org/10.1007/S11023-024-09685-X
- Purificato et al. (2025) Erasmo Purificato, Hannan Javed Mahadik, Ludovico Boratto, and Ernesto William De Luca. 2025. GNN’s FAME: Fairness-Aware MEssages for Graph Neural Networks. In Proceedings of the 33rd ACM Conference on User Modeling, Adaptation and Personalization, UMAP 2025, New York City, NY, USA, June 16-19, 2025. ACM, 301–306. https://doi.org/10.1145/3699682.3728324
@@
- Qin et al. (2024) Yifang Qin, Hongjun Wu, Wei Ju, Xiao Luo, and Ming Zhang. 2024. A Diffusion Model for POI Recommendation. ACM Trans. Inf. Syst. 42, 2 (2024), 54:1–54:27. https://doi.org/10.1145/3624475
@@
- Rendle et al. (2009) Steffen Rendle, Christoph Freudenthaler, Zeno Gantner, and Lars Schmidt-Thieme. 2009. BPR: Bayesian Personalized Ranking from Implicit Feedback. In UAI 2009, Proceedings of the Twenty-Fifth Conference on Uncertainty in Artificial Intelligence, Montreal, QC, Canada, June 18-21, 2009, Jeff A. Bilmes and Andrew Y. Ng (Eds.). AUAI Press, 452–461. https://www.auai.org/uai2009/papers/UAI2009_0139_48141db02b9f0b02bc7158819ebfa2c7.pdf
@@
- Rombach et al. (2022) Robin Rombach, Andreas Blattmann, Dominik Lorenz, Patrick Esser, and Björn Ommer. 2022. High-Resolution Image Synthesis with Latent Diffusion Models. In IEEE/CVF Conference on Computer Vision and Pattern Recognition, CVPR 2022, New Orleans, LA, USA, June 18-24, 2022. IEEE, 10674–10685. https://doi.org/10.1109/CVPR52688.2022.01042
@@
- Sarwar et al. (2001) Badrul Munir Sarwar, George Karypis, Joseph A. Konstan, and John Riedl. 2001. Item-based collaborative filtering recommendation algorithms. In Proceedings of the Tenth International World Wide Web Conference, WWW 10, Hong Kong, China, May 1-5, 2001, Vincent Y. Shen, Nobuo Saito, Michael R. Lyu, and Mary Ellen Zurko (Eds.). ACM, 285–295. https://doi.org/10.1145/371920.372071
@@
- Shen et al. (2024) Xudong Shen, Chao Du, Tianyu Pang, Min Lin, Yongkang Wong, and Mohan S. Kankanhalli. 2024. Finetuning Text-to-Image Diffusion Models for Fairness. In The Twelfth International Conference on Learning Representations, ICLR 2024, Vienna, Austria, May 7-11, 2024. OpenReview.net. https://openreview.net/forum?id=hnrB5YHoYu
@@
- Shenbin et al. (2020) Ilya Shenbin, Anton Alekseev, Elena Tutubalina, Valentin Malykh, and Sergey I. Nikolenko. 2020. RecVAE: A New Variational Autoencoder for Top-N Recommendations with Implicit Feedback. In WSDM ’20: The Thirteenth ACM International Conference on Web Search and Data Mining, Houston, TX, USA, February 3-7, 2020, James Caverlee, Xia (Ben) Hu, Mounia Lalmas, and Wei Wang (Eds.). ACM, 528–536. https://doi.org/10.1145/3336191.3371831
- Steck (2019) Harald Steck. 2019. Embarrassingly Shallow Autoencoders for Sparse Data. In The World Wide Web Conference, WWW 2019, San Francisco, CA, USA, May 13-17, 2019, Ling Liu, Ryen W. White, Amin Mantrach, Fabrizio Silvestri, Julian J. McAuley, Ricardo Baeza-Yates, and Leila Zia (Eds.). ACM, 3251–3257. https://doi.org/10.1145/3308558.3313710
@@
- Sun et al. (2020) Jianing Sun, Wei Guo, Dengcheng Zhang, Yingxue Zhang, Florence Regol, Yaochen Hu, Huifeng Guo, Ruiming Tang, Han Yuan, Xiuqiang He, and Mark Coates. 2020. A Framework for Recommending Accurate and Diverse Items Using Bayesian Graph Convolutional Neural Networks. In KDD ’20: The 26th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, Virtual Event, CA, USA, August 23-27, 2020, Rajesh Gupta, Yan Liu, Jiliang Tang, and B. Aditya Prakash (Eds.). ACM, 2030–2039. https://doi.org/10.1145/3394486.3403254
@@
- Togashi et al. (2024) Riku Togashi, Kenshi Abe, and Yuta Saito. 2024. Scalable and Provably Fair Exposure Control for Large-Scale Recommender Systems. In Proceedings of the ACM on Web Conference 2024, WWW 2024, Singapore, May 13-17, 2024, Tat-Seng Chua, Chong-Wah Ngo, Ravi Kumar, Hady W. Lauw, and Roy Ka-Wei Lee (Eds.). ACM, 3307–3318. https://doi.org/10.1145/3589334.3645390
@@
- Vassøy and Langseth (2024) Bjørnar Vassøy and Helge Langseth. 2024. Consumer-side fairness in recommender systems: a systematic survey of methods and evaluation. Artif. Intell. Rev. 57, 4 (2024), 101. https://doi.org/10.1007/S10462-023-10663-5
- Vignac et al. (2023) Clément Vignac, Igor Krawczuk, Antoine Siraudin, Bohan Wang, Volkan Cevher, and Pascal Frossard. 2023. DiGress: Discrete Denoising diffusion for graph generation. In The Eleventh International Conference on Learning Representations, ICLR 2023, Kigali, Rwanda, May 1-5, 2023. OpenReview.net. https://openreview.net/forum?id=UaAD-Nu86WX
- Walker et al. (2022) Joojo Walker, Ting Zhong, Fengli Zhang, Qiang Gao, and Fan Zhou. 2022. Recommendation via Collaborative Diffusion Generative Model. In Knowledge Science, Engineering and Management - 15th International Conference, KSEM 2022, Singapore, August 6-8, 2022, Proceedings, Part III (Lecture Notes in Computer Science, Vol. 13370), Gérard Memmi, Baijian Yang, Linghe Kong, Tianwei Zhang, and Meikang Qiu (Eds.). Springer, 593–605. https://doi.org/10.1007/978-3-031-10989-8_47
@@
- Wang et al. (2017) Jun Wang, Lantao Yu, Weinan Zhang, Yu Gong, Yinghui Xu, Benyou Wang, Peng Zhang, and Dell Zhang. 2017. IRGAN: A Minimax Game for Unifying Generative and Discriminative Information Retrieval Models. In Proceedings of the 40th International ACM SIGIR Conference on Research and Development in Information Retrieval, Shinjuku, Tokyo, Japan, August 7-11, 2017, Noriko Kando, Tetsuya Sakai, Hideo Joho, Hang Li, Arjen P. de Vries, and Ryen W. White (Eds.). ACM, 515–524. https://doi.org/10.1145/3077136.3080786
- Wang et al. (2023) Wenjie Wang, Yiyan Xu, Fuli Feng, Xinyu Lin, Xiangnan He, and Tat-Seng Chua. 2023. Diffusion Recommender Model. In Proceedings of the 46th International ACM SIGIR Conference on Research and Development in Information Retrieval, SIGIR 2023, Taipei, Taiwan, July 23-27, 2023, Hsin-Hsi Chen, Wei-Jou (Edward) Duh, Hen-Hsen Huang, Makoto P. Kato, Josiane Mothe, and Barbara Poblete (Eds.). ACM, 832–841. https://doi.org/10.1145/3539618.3591663
@@
- Wu et al. (2024) Likang Wu, Zhi Zheng, Zhaopeng Qiu, Hao Wang, Hongchao Gu, Tingjia Shen, Chuan Qin, Chen Zhu, Hengshu Zhu, Qi Liu, Hui Xiong, and Enhong Chen. 2024. A survey on large language models for recommendation. World Wide Web (WWW) 27, 5 (2024), 60. https://doi.org/10.1007/S11280-024-01291-2
@@
- Wu et al. (2023) Shiwen Wu, Fei Sun, Wentao Zhang, Xu Xie, and Bin Cui. 2023. Graph Neural Networks in Recommender Systems: A Survey. ACM Comput. Surv. 55, 5 (2023), 97:1–97:37. https://doi.org/10.1145/3535101
@@
- Xu et al. (2022) Minkai Xu, Lantao Yu, Yang Song, Chence Shi, Stefano Ermon, and Jian Tang. 2022. GeoDiff: A Geometric Diffusion Model for Molecular Conformation Generation. In The Tenth International Conference on Learning Representations, ICLR 2022, Virtual Event, April 25-29, 2022. OpenReview.net. https://openreview.net/forum?id=PzcvxEMzvQC
@@
- Yang et al. (2016) Dingqi Yang, Daqing Zhang, and Bingqing Qu. 2016. Participatory Cultural Mapping Based on Collective Behavior Data in Location-Based Social Networks. ACM Trans. Intell. Syst. Technol. 7, 3 (2016), 30:1–30:23. https://doi.org/10.1145/2814575
@@
- Yang et al. (2024) Ling Yang, Zhilong Zhang, Yang Song, Shenda Hong, Runsheng Xu, Yue Zhao, Wentao Zhang, Bin Cui, and Ming-Hsuan Yang. 2024. Diffusion Models: A Comprehensive Survey of Methods and Applications. ACM Comput. Surv. 56, 4 (2024), 105:1–105:39. https://doi.org/10.1145/3626235
@@
- Yu et al. (2024) Junliang Yu, Xin Xia, Tong Chen, Lizhen Cui, Nguyen Quoc Viet Hung, and Hongzhi Yin. 2024. XSimGCL: Towards Extremely Simple Graph Contrastive Learning for Recommendation. IEEE Trans. Knowl. Data Eng. 36, 2 (2024), 913–926. https://doi.org/10.1109/TKDE.2023.3288135
@@
- Zhao et al. (2022b) Minghao Zhao, Le Wu, Yile Liang, Lei Chen, Jian Zhang, Qilin Deng, Kai Wang, Xudong Shen, Tangjie Lv, and Runze Wu. 2022b. Investigating Accuracy-Novelty Performance for Graph-based Collaborative Filtering. In SIGIR ’22: The 45th International ACM SIGIR Conference on Research and Development in Information Retrieval, Madrid, Spain, July 11 - 15, 2022, Enrique Amigó, Pablo Castells, Julio Gonzalo, Ben Carterette, J. Shane Culpepper, and Gabriella Kazai (Eds.). ACM, 50–59. https://doi.org/10.1145/3477495.3532005
@@
- Zhao et al. (2022a) Wayne Xin Zhao, Yupeng Hou, Xingyu Pan, Chen Yang, Zeyu Zhang, Zihan Lin, Jingsen Zhang, Shuqing Bian, Jiakai Tang, Wenqi Sun, Yushuo Chen, Lanling Xu, Gaowei Zhang, Zhen Tian, Changxin Tian, Shanlei Mu, Xinyan Fan, Xu Chen, and Ji-Rong Wen. 2022a. RecBole 2.0: Towards a More Up-to-Date Recommendation Library. In Proceedings of the 31st ACM International Conference on Information & Knowledge Management, Atlanta, GA, USA, October 17-21, 2022, Mohammad Al Hasan and Li Xiong (Eds.). ACM, 4722–4726. https://doi.org/10.1145/3511808.3557680